### v1

@@ 1 Introduction
Here, we consider a planar jet, i.e., the flow is injected through a plane slit of half-width $h$ in a computational box periodic in the vertical direction $z$. We adopt a shear-thinning fluid in which the viscosity decreases non-linearly for increasing values of the shear rate. This setup is representative of the outflow of a river into the sea or the ocean, in which a stream of freshwater rich in bacteria or microalgae is delivered into sea water. Indeed, the presence of bacteria and microalgae in suspension has been shown to introduce non-Newtonian features into the fluid, such as shear-dependent viscosity (Al-Asheh et al., 2002; Zhang et al., 2013). Differences in salinity, temperature and density between the freshwater stream and the salt water can impede mixing, thus leading to the formation of a stratified flow with a (thin) layer of freshwater flowing over salt water. The freshwater layer rich in microalgae and bacteria is reminiscent of our computational setup, where the height of the domain (the thickness of the freshwater layer) is much smaller than the other two dimensions.
@@
2 Setup and method
2.1 Problem setup
We have addressed this study by means of three-dimensional direct numerical simulations. The planar jet is generated by fluid injected with a uniform velocity $U$ through a planar slit spanning the entire height of the domain, and it develops over a stream-wise distance $L_{x}=160h$ in a computational box of width $L_{y}=240h$ and height $L_{z}$. The vertical length $L_{z}$ is varied among simulations: we consider five distinct simulations with $L_{z}=0.83h,1.67h,3.33h,6.67h,13.33h$, respectively. The thinnest domain ($L_{z}=0.83h$) introduces a strong vertical constraint in order to allow the development of a two-dimensional flow, which is progressively relaxed as $L_{z}$ is increased while maintaining a thin computational domain, so $L_{z}\ll L_{x}\sim L_{y}$.
The flow dynamics are simulated by solving the Navier-Stokes equations:
$\nabla\cdot\boldsymbol{u}=0,$ (1)
@@ 2.1 Problem setup
$\rho\left(\frac{\partial\mathbf{u}}{\partial t}+\mathbf{u}\cdot\nabla{\mathbf{u}}\right)=-\nabla p+\nabla\left[\mu\left(\nabla\mathbf{u}+\nabla\mathbf{u}^{T}\right)\right],$ (2)
@@
where $\mathbf{u}$ is the local flow velocity, $\rho$ the density, $p$ the pressure and $\mu$ the local viscosity. We adopt an inelastic, shear-thinning fluid, whose behaviour is defined via the Carreau fluid model (Bird et al., 1974). The local viscosity $\mu$ depends on the local shear rate $\dot{\gamma}$ as:
$\mu(\dot{\gamma})=\mu_{\infty}+\left(\mu_{0}+\mu_{\infty}\right)\left[1+\left(\lambda\dot{\gamma}\right)^{2}\right]^{\frac{n-1}{2}},$ (3)
where $\lambda$ is the fluid consistency index, and $\mu_{0}$ and $\mu_{\infty}$ are the zero-shear viscosity and the viscosity for $\dot{\gamma}\rightarrow\infty$, respectively. We set the power-law index $n=0.2$, thereby obtaining a strong shear-thinning effect. We select a low value of the inlet Reynolds number (ratio of inertial to viscous effects), $Re=\rho hU/\mu_{0}=20$. It should be noted that Newtonian planar jets are laminar at this value of $Re$ (Deo et al., 2008; Sureshkumar & Beris, 1995; Sato & Sakao, 1964; Soligo & Rosti, 2023), thus any turbulent motion is caused exclusively by the shear-thinning in the flow. Turbulence however is still Newtonian, as it originates by the prevalence of inertial over viscous terms. The non-Newtonian character of the flow promotes indeed the onset of the instability, so the transition to turbulence is at markedly lower $Re$ compared to Newtonian planar jets (Ray & Zaki, 2015; Yamani et al., 2023; Soligo & Rosti, 2023). The non-Newtonian contribution is described using the Carreau number, defined as $Cu=h\lambda/U=100$.
Our numerical setup is easily testable by experiments (recent experimental work by Yamani et al. (2023) addressed the flow of a viscoelastic planar jet at low Reynolds), whereas the effect of physical confinement is attenuated by the shear-thinning characteristic of the fluid, which reduces the extent of the shear layer (viscosity decreases at the wall boundaries in the experimental setup).
2.2 Numerical method
@@
The motion of the incompressible, shear-thinning fluid is solved by simulating the Navier-Stokes equations with variable viscosity using the in-house solver Fujin (https://groups.oist.jp/cffu/code). The Navier-Stokes equations are discretized on a uniform, staggered, Cartesian grid; the fluid velocities are located at the cell faces, whereas pressure and viscosity at the cell centers. The fluid viscosity is updated at every time-step following Eq. 3. The spatial derivatives are approximated using second-order finite differences in all directions. The system is advanced in time through a second-order Adams-Bashforth scheme coupled with a fractional step method (Kim & Moin, 1985). The divergence-free velocity field is enforced by solving the Poisson equation for the pressure using an efficient solver based on the fast Fourier transform. We resort to the domain decomposition library 2decomp (http://www.2decomp.org) and the MPI protocol to parallelize the solver.
We adopt a uniform grid spacing in all spatial directions, in $x$ and $y$ the domain is discretized using $N_{x}\times N_{y}=1536\times 2304$ grid points, while the number of points in the $z$ direction depends on the height of the domain, namely $N_{z}=8,16,32,64,128$ for increasing heights. The jet fluid is issued through a slit at $x=0$; at this boundary we impose no-slip and no-penetration boundary conditions, exception made for the inlet part, where we impose a plug-flow velocity profile. At the outlet boundary ($x=L_{x}$) we use a non-reflective boundary condition (Orlanski, 1976). At the side boundaries ($y=0$ and $y=L_{y}$) we impose free-slip and no penetration boundary conditions. Lastly, at the top and bottom boundaries ($z=0$ and $z=L_{z}$) we impose periodic boundary conditions.
@@ 3 Results
Next, we inspect the energy spectra in the bottom panel of Fig. 1. Note that we compute the velocity spectra in time rather than in space by recording velocity data over time from a probe placed on the centerline of the jet, similarly to what done in experiments. Computing the power spectra in time rather than in the vertical direction allows us to have a wider energy spectrum that is not limited by the height of the domain. The equivalence of time and space spectra has been demonstrated in the past (Namer & Ötügen, 1988; Soligo & Rosti, 2023). As clearly shown in the figure, $L_{z}$ influences the energy cascade, which depicts a different behavior depending on the regime of turbulence. First, the least constrained jet shows the conventional $f^{-5/3}$ scaling for three-dimensional turbulence (Kolmogorov, 1941). Consistently, we observe the energy cascade typical of three-dimensional turbulence: the jet instability gives energy to the flow and it generates large structures that break down in progressively smaller and smaller eddies. As the vortices move downstream, the characteristic shear rate reduces – hence viscosity increases – and energy is dissipated. We expect to recover the power-law scaling for the three-dimensional turbulence energy cascade as the turbulent motions are Newtonian: they are generated by the competition of inertial and viscous terms (Soligo & Rosti, 2023). Eventually, dissipation becomes relevant at every scale and the cascade is impeded: the scaling $f^{-5/3}$ is not present at $x=120h$. On the other hand, the most constrained case exhibit two-dimensional flow and a scaling of $f^{-3}$. Here, the three-dimensional cascade is clearly disrupted and two-dimensional phenomena become dominant (Kraichnan, 1967; Batchelor, 1969). The change in the flowing regime observed here is due only to the vertical confinement: we adopt a non-Newtonian fluid model which is characterized by the presence of shear-thinning alone (there are no viscoelastic effects). The spectrum becomes steeper as $x$ increases and it eventually seems to saturate at $f^{-5}$. The steepening of the energy spectrum agrees with the appearance of dispersed large-size coherent vortices in the flow (Basdevant et al., 1981; McWilliams, 1984; Benzi et al., 1986; Legras et al., 1988). Note that the change in the vertical constraint also alters the instability in the region close to the inlet: for strong vertical constraints (small $L_{z}$) we observe a flapping motion of the shear layers, whereas puffing motion dominates when the constraint is relaxed (large $L_{z}$). The flapping dynamic is associated with the antisymmetric, sinuous mode, that destabilizes the flow more substantially, thus injecting more energy (Mattingly & Criminale Jr, 1971), as also observed by the energy spectra close to the inlet which are shifted upwards in 2D compared to the 3D case. At last, the mixed-dimensional planar jet exhibits features from both three- and two- dimensional planar jets: the scaling $f^{-5/3}$ is found close to the inlet and it changes towards $f^{-3}$ downstream. The spectrum becomes slightly steeper further downstream, consistent with the appearance of the coherent vortices. Similarly to the two-dimensional jet, the flow is more energetic close to the inlet. Very close to the inlet, the vorticity fluctuations look closely related to those in the three-dimensional case, implying the existence of puffing events connected to the varicose mode (Sato, 1960; Mattingly & Criminale Jr, 1971). However, flapping motions become soon dominant and the macroscopic vorticity structures resemble those from the two-dimensional planar jet, thus indicating the presence of the more energetic sinuous mode.
@@
Figure 2: Third-order structure functions $S_{3}$ of the longitudinal velocity fluctuations for different vertical constraints $L_{z}$. We compute $S_{3}$ at $x=40h$ at the flow centerline for each planar jet: $L_{z}=0.83h,1.67h,3.33h,6.67h,13.33h$. We show the absolute value $\left|S_{3}\right|$ for clarity, and we indicate with colors whether $S_{3}$ is positive (blue) or negative (red), so that the corresponding turbulent scales are either two- or three- dimensional. Consequently, the scaling for $\left|S_{3}\right|$ changes as $L_{z}$ is shifted towards larger values. We identify two scalings: $S_{3}\sim r^{3}$ if turbulence is two-dimensional, corresponding to a direct cascade of enstrophy, and $S_{3}\sim-r$ if it is three-dimensional, thus denoting a direct cascade of energy. The plots are shifted vertically for better readability.
@@
To better characterise the different nature of the turbulent fluctuations at each scale, we now calculate the longitudinal velocity differences $\Delta u(r)=\left(\boldsymbol{u}\left(\boldsymbol{x}+\boldsymbol{r}\right)-\boldsymbol{u}\left(\boldsymbol{x}\right)\right)\cdot\boldsymbol{r}/\left|\boldsymbol{r}\right|$. Concretely, we introduce the third-order structure function $S_{3}(r)=\langle\left(\Delta u\right)^{3}\rangle$, where the angle brackets indicate averaging in time and in space, shown in Fig. 2 for the different cases analysed. A remarkable property of $S_{3}$ is that it denotes whether the flow scales are two- or three- dimensional depending on its sign (Lindborg, 1999; Kolmogorov, 1991), and it can help understanding the direction of the energy and enstrophy fluxes (Cho & Lindborg, 2001; Bernard, 1999; Cerbus & Chakraborty, 2017). In the particular case of two-dimensional turbulence, $S_{3}$ is positive (Lindborg, 1999) whereas it is negative if turbulence is three-dimensional (Kolmogorov, 1991). Indeed, we observe that $S_{3}$ is positive (two-dimensional flow) for the most constrained case ($L_{z}=0.83h$) whereas is is negative (three-dimensional flow) for the least constrained case ($L_{z}=6.67h,13.33h$). Furthermore, the scaling of $S_{3}$ outlines the preferred cascade process, thus indicating the direct cascade of enstrophy in the two-dimensional case ($S_{3}\sim r^{3}$) and the direct cascade of energy in the three-dimensional counterpart ($S_{3}\sim-r$). We do not find evidence of an inverse mechanism of energy transfer toward the large scales in the two-dimensional case. This is not surprising if we consider that energy is injected in the flow through the planar slit, thus not forcing the flow at any intermediate scale (Boffetta, 2007; Boffetta & Ecke, 2012). More interestingly, turbulence is characterized by a mixed-dimensional regime for the right choice of $L_{z}$. Two- and three- dimensional scales are present simultaneously in the flow for $L_{z}=1.67h,3.33h$. In these cases, large scales are two-dimensional while the small ones are three-dimensional, with the transition between regimes happening at some intermediate scale. This transition is strongly dependent on the height of the domain and it can be delayed further in $x$ for increasing values of $L_{z}$.
@@
Figure 3: Third-order structure functions $S_{3}$ of the longitudinal velocity fluctuations for different distances from the inlet. Panel a) shows $S_{3}$ computed at distances $x$ uniformly separated along the jet centerline ($x=20h,30h,\ldots,140h$) for the mixed-dimensional planar jet with $L_{z}=3.33h$. We indicate whether the turbulent scales are two- or three- dimensional with colors, similarly to Fig. 2. The plots are shifted vertically for better readability. The inset shows the scale $r$ at which $S_{3}$ changes its sign. Red and blue markers are used for fully $3$D and $2$D structure functions respectively, while purple markers for those showing both regimes. Note that, if $S_{3}$ does not change sign, we mark either the largest three-dimensional scale or the smallest two-dimensional one. Panel b) depicts $S_{3}$ for the least constrained ($L_{z}=13.33h$) or $3$D jet, and panel c) for the most constrained ($L_{z}=0.83h$) or $2$D jet. $S_{3}$ is plotted in a similar fashion than a), where we computed $S_{3}$ at three distances $x=40h,80h,120h$ at the jet centerline. In these cases, $S_{3}$ has the same sign throughout $x$, thus observing a single scaling at all distances dependent on the dimensionality of the flow.
@@
To further investigate the transition between regimes, we choose a mixed-dimensional planar jet, in particular $L_{z}=3.33h$, and we calculate repeatedly $S_{3}$ at several distances from the inlet. Results are summarized in Fig. 3a. Close to the inlet the flow is three-dimensional: the effect of the vertical constraint does not hinder the development of the three-dimensional regime. The largest flow scale is smaller than the geometrical constraint, thus not compromising the energy cascade of the three-dimensional flow. We report the presence of a single scaling, $S_{3}\sim-r$, typical of the direct energy cascade. At intermediate distances, $x=\left[40h,100h\right]$, a mixed-dimensional regime appears: both two- and three-dimensional regimes coexist, with three-dimensional turbulence characterizing the small scales and two-dimensional turbulence characterizing instead the large scales. Here, the vertical confinement hinders the development of three-dimensional turbulence at the largest scales, while it has no effect on the three dimensional turbulence at the smallest scales. In this case, we have the simultaneous presence of a direct energy cascade ($S_{3}\sim-r$) at small scales and a direct enstrophy cascade ($S_{3}\sim r^{3}$) at large scales. We report in the inset of Fig. 3a the scale at which the structure function changes sign, i.e., the scale where the flow transitions from $3$D to $2$D. We observe that, as soon as the jet thickness reaches the size of the vertical confinement, the largest scales become two-dimensional. However, some scales between the largest ones and the vertical confinement remain three-dimensional, causing the largest $3$D structures to be anisotropic. Moving downstream, the anisotropy of the largest $3$D structures grows till a value of around $3L_{z}$ at $x\approx 60h$, after which the flow becomes more and more two-dimensional. Eventually, we observe that the flow becomes completely two-dimensional at the farthest distances from the inlet, $x\geq 120h$, where the characteristic flow scales are the largest and the vertical confinement impedes the development of three-dimensional flow at all scales. We indeed observe a scaling of the structure function $S_{3}$ compatible with the direct enstrophy cascade ($S_{3}\sim r^{3}$). The range of scales observed at each distance from the inlet depends on two factors: the characteristic length scale of the jet (the jet thickness), and the local viscosity. Both of these quantities increase with increasing distance from the inlet of the jet; the jet thickness determines the largest scale in the flow, while the local viscosity is among the factors determining the eventual development of three-dimensional turbulent motions.

### v4

@@ 1 Introduction
Here, we consider a planar jet, i.e., the flow is injected through a plane slit of half-width $h$ in a computational box periodic in the vertical direction $z$. We adopt a shear-thinning fluid in which the viscosity decreases non-linearly for increasing values of the shear rate. A similar flow configuration, although at a much higher characteristic Reynolds number, can be found at the outflow of a river into the sea. Differences in salinity, temperature and density between the freshwater stream and the salt water can impede mixing, thus leading to the formation of a stratified flow with a (thin) layer of freshwater flowing over salt water. The presence of suspended bacteria and microalgae in the freshwater stream grants non-Newtonian features to the fluid, such as shear-dependent viscosity (Al-Asheh et al., 2002; Zhang et al., 2013). Our simulations are performed at a much lower value of the Reynolds number, hence there is no direct application of our findings to the flowing regimes found at the outflow of a river; the low value of the Reynolds number allows instead for direct comparisons with experiments. The present configuration can be easily achieved in laboratory experiments; the shear rheology of the fluid used in our simulations corresponds to that of a 100:60 mM CPyCl:NaSal wormlike micelle solution (Haward et al., 2021). This work thus constitutes a preliminary step in understanding the flow dynamics of geophysical flows, such as the outflow of a river rich in bacteria or microalgae, which are instead characterized by a much larger Reynolds number. The selected parameters make our numerical setup easily testable by experiments (recent experimental work by Yamani et al. (2023) addressed the flow of a viscoelastic planar jet at low Reynolds), whereas the effect of physical confinement is attenuated by the shear-thinning characteristic of the fluid, which reduces the extent of the shear layer (viscosity decreases at the wall boundaries in the experimental setup).
@@
2 Numerical method
The motion of the incompressible, shear-thinning fluid is governed by the mass, Eq. 1, and momentum, Eq. 2, conservation equations.
$\nabla\cdot\boldsymbol{u}=0$ (1)
@@ 2 Numerical method
$\rho\left(\frac{\partial\mathbf{u}}{\partial t}+\mathbf{u}\cdot\nabla{\mathbf{u}}\right)=-\nabla p+\nabla\cdot\left[\mu(\dot{\gamma})\left(\nabla\mathbf{u}+\nabla\mathbf{u}^{T}\right)\right]$ (2)
@@
In the above equations, $\mathbf{u}$ is the local flow velocity, $\rho$ the density, $p$ the pressure and $\mu$ the local viscosity. We adopt an inelastic, shear-thinning fluid, whose behaviour is defined via the Carreau fluid model (Bird et al., 1974). The local viscosity $\mu$ depends on the local shear rate $\dot{\gamma}$ as:
$\mu(\dot{\gamma})=\mu_{\infty}+\left(\mu_{0}-\mu_{\infty}\right)\left[1+\left(\lambda\dot{\gamma}\right)^{2}\right]^{\frac{n-1}{2}},$ (3)
[…]
@@
The Navier-Stokes equations are discretized on a uniform, staggered, Cartesian grid; the fluid velocities are located at the cell faces, whereas pressure and viscosity are defined at the cell centers. The fluid viscosity is updated at every time-step following Eq. 3. The spatial derivatives are approximated using second-order finite differences in all directions. The system is advanced in time through a second-order Adams-Bashforth scheme coupled with a fractional step method (Kim & Moin, 1985). The divergence-free velocity field is enforced by solving the Poisson equation for the pressure using an efficient solver based on the fast Fourier transform. We resort to the domain decomposition library 2decomp (http://www.2decomp.org) and the MPI protocol to parallelize the solver. The numerical solver is implemented in the in-house solver Fujin (https://groups.oist.jp/cffu/code).
3 Problem setup
[…]
We select a low value of the inlet Reynolds number (ratio of inertial to viscous effects), $Re=\rho hU/\mu_{0}=20$. It should be noted that Newtonian planar jets are laminar at this value of $Re$ (Deo et al., 2008; Sureshkumar & Beris, 1995; Sato & Sakao, 1964; Soligo & Rosti, 2023), thus any turbulent motion is caused exclusively by the shear-thinning in the flow. Turbulence however is still Newtonian, as it originates by the prevalence of inertial over viscous terms. The non-Newtonian character of the flow promotes indeed the onset of the instability, so the transition to turbulence is at markedly lower $Re$ compared to Newtonian planar jets (Ray & Zaki, 2015; Yamani et al., 2023; Soligo & Rosti, 2023). The non-Newtonian contribution is described using the Carreau number, defined as $Cu=h\lambda/U=100$. The ratio between the zero-shear viscosity and the infinite-shear viscosity is set to $\mu_{0}/\mu_{\infty}=50$.
To verify the independence of our results on the specific parameters selected, we perform two additional simulations at a set $L_{z}=3.33h$ and we double either the Reynolds ($Re=40$) or the Carreau ($Cu=200$) number. The reference case ($L_{z}=3.33h$, $Cu=100$ and $Re=20$) exhibits a mixed-dimensional turbulent regime, where features from two- and three-dimensional turbulence are found simultaneously in the flow (see Fig. 1). In the cases at higher $Re$ or $Cu$, we expect turbulence to be enhanced. The Reynolds number is increased by reducing the zero-shear viscosity (and consequently the infinite-shear viscosity, defined as $\mu_{\infty}=\mu_{0}/50$), whereas the Carreau number is increased by changing the fluid consistency index $\lambda$, keeping the zero-shear and infinite shear viscosities unchanged. This way, the transition towards the infinite-shear viscosity occurs at a smaller shear rate compared to the case at $Cu=100$.
We adopt a uniform grid spacing in all spatial directions for all cases; in $x$ and $y$ the domain is discretized using $N_{x}\times N_{y}=1536\times 2304$ grid points, while the number of points in the $z$ direction depends on the height of the domain, namely $N_{z}=8,16,32,64,128$ for increasing heights. We ensured that the grid resolution is adequate by computing the ratio between the grid spacing $\Delta$ and the mean Kolmogorov length scale $\eta$:
$\eta=\left(\frac{\langle\nu\rangle^{3}}{\langle\varepsilon\rangle}\right)^{1/4},$ (4)
where $\nu$ is the local kinematic viscosity and $\varepsilon$ is the viscous dissipation; angle brackets indicate averaging in time and in the vertical direction $z$. In the cases characterized by the highest turbulence intensity, $Re=40$ and $Cu=200$, the Kolmogorov length scale is always $\eta\gtrsim 0.5\Delta$ (where $\Delta$ is the grid spacing, uniform in the three directions). The smallest values of the Kolmogorov length scale are encountered at the jet centerline within the region $15h<x<30h$; beyond $x=40h$ the Kolmogorov length scale is always larger than the grid spacing, thus ensuring that the grid resolution chosen is adequate for all cases.
@@ 4.1 Effect of vertical confinement
Next, we inspect the energy spectra in the bottom panel of Fig. 1. Note that we compute the velocity spectra in time rather than in space by recording velocity data over time from a probe placed on the centerline of the jet, similarly to what done in experiments. Computing the power spectra in time rather than in the vertical direction allows us to have a wider energy spectrum that is not limited by the height of the domain. The equivalence of time and space spectra has been demonstrated in the past (Namer & Ötügen, 1988; Soligo & Rosti, 2023). As clearly shown in the figure, $L_{z}$ influences the energy cascade, which depicts a different behavior depending on the regime of turbulence. First, the least constrained jet shows the conventional $f^{-5/3}$ scaling for three-dimensional turbulence (Kolmogorov, 1941). Consistently, we observe the energy cascade typical of three-dimensional turbulence: the jet instability gives energy to the flow and it generates large structures that break down in progressively smaller and smaller eddies. As the vortices move downstream, the characteristic shear rate reduces – hence viscosity increases – and energy is dissipated. We expect to recover the power-law scaling for the three-dimensional turbulence energy cascade as the turbulent motions are Newtonian: they are generated by the competition of inertial and viscous terms (Soligo & Rosti, 2023). Eventually, dissipation becomes relevant at every scale and the cascade is impeded: the scaling $f^{-5/3}$ is not present at $x=120h$. On the other hand, the most constrained case exhibits two-dimensional flow and a scaling of $f^{-3}$. Here, the three-dimensional cascade is clearly disrupted and two-dimensional phenomena become dominant (Kraichnan, 1967; Batchelor, 1969). The change in the flowing regime observed here is due only to the vertical confinement: we adopt a non-Newtonian fluid model which is characterized by the presence of shear-thinning alone (there are no viscoelastic effects). The spectrum becomes steeper as $x$ increases and it eventually seems to saturate at $f^{-5}$. The steepening of the energy spectrum agrees with the appearance of dispersed large-size coherent vortices in the flow (Basdevant et al., 1981; McWilliams, 1984; Benzi et al., 1986; Legras et al., 1988). Note that the change in the vertical constraint also alters the instability in the region close to the inlet: for strong vertical constraint (small $L_{z}$) we observe a flapping motion of the shear layers, whereas puffing motion dominates when the constraint is relaxed (large $L_{z}$). The flapping dynamic is associated with the antisymmetric, sinuous mode, that destabilizes the flow more substantially, thus injecting more energy (Mattingly & Criminale Jr, 1971), as also observed by the energy spectra close to the inlet, which are shifted upwards in 2D compared to the 3D case. At last, the mixed-dimensional planar jet exhibits features from both three- and two- dimensional planar jets: the scaling $f^{-5/3}$ is found close to the inlet and it changes towards $f^{-3}$ downstream. The spectrum becomes slightly steeper further downstream, consistent with the appearance of the coherent vortices. Similarly to the two-dimensional jet, the flow is more energetic close to the inlet. Very close to the inlet, the vorticity fluctuations look closely related to those in the three-dimensional case, implying the existence of puffing events connected to the varicose mode (Sato, 1960; Mattingly & Criminale Jr, 1971). However, flapping motions become soon dominant and the macroscopic vorticity structures resemble those from the two-dimensional planar jet, thus indicating the presence of the more energetic sinuous mode.
@@
Figure 2: Third-order structure functions $S_{3}$ of the longitudinal velocity fluctuations for different vertical constraints $L_{z}$. We compute $S_{3}$ at $x=40h$ at the flow centerline for each planar jet: $L_{z}=0.83h,1.67h,3.33h,6.67h,13.33h$. We show the absolute value $\left|S_{3}\right|$ for clarity, and we indicate with colors whether $S_{3}$ is positive (blue) or negative (red), so that the corresponding turbulent scales are either two- or three- dimensional. Consequently, the scaling for $\left|S_{3}\right|$ changes as $L_{z}$ is shifted towards larger values. We report also the two scalings: $S_{3}\sim r^{3}$ if turbulence is two-dimensional, corresponding to a direct cascade of enstrophy, and $S_{3}\sim-r$ if it is three-dimensional, thus denoting a direct cascade of energy. The plots are shifted vertically for better readability.
@@
To better characterise the different nature of the turbulent fluctuations at each scale, we now calculate the longitudinal velocity differences $\Delta u(r)=\left(\boldsymbol{u}\left(\boldsymbol{x}+\boldsymbol{r}\right)-\boldsymbol{u}\left(\boldsymbol{x}\right)\right)\cdot\boldsymbol{r}/\left|\boldsymbol{r}\right|$. Concretely, we introduce the third-order structure function $S_{3}(r)=\langle\left(\Delta u\right)^{3}\rangle$, where the angle brackets indicate averaging in time and in space, shown in Fig. 2 for the different cases analysed. Appendix A reports in detail how the computation of the structure function was performed. A remarkable property of $S_{3}$ is that it denotes whether the flow scales are two- or three- dimensional depending on its sign (Lindborg, 1999; Kolmogorov, 1991), and it can help understanding the direction of the energy and enstrophy fluxes (Cho & Lindborg, 2001; Bernard, 1999; Cerbus & Chakraborty, 2017). In the particular case of two-dimensional turbulence, $S_{3}$ is positive (Lindborg, 1999) whereas it is negative if turbulence is three-dimensional (Kolmogorov, 1991). Indeed, we observe that $S_{3}$ is positive (two-dimensional flow) for the most constrained case ($L_{z}=0.83h$) whereas it is negative (three-dimensional flow) for the least constrained case ($L_{z}=6.67h,13.33h$). Furthermore, the scaling of $S_{3}$ outlines the preferred cascade process, thus indicating the direct cascade of enstrophy in the two-dimensional case ($S_{3}\sim r^{3}$) and the direct cascade of energy in the three-dimensional counterpart ($S_{3}\sim-r$). While the scaling $S_{3}\sim r^{3}$ is visible for two-dimensional turbulence, we do not observe a clear $S_{3}\sim-r$ scaling for all three-dimensional turbulence cases. We do not find evidence of an inverse mechanism of energy transfer toward the large scales in the two-dimensional case. This is not surprising if we consider that energy is injected in the flow through the planar slit, thus not forcing the flow at any intermediate scale (Boffetta, 2007; Boffetta & Ecke, 2012). More interestingly, turbulence is characterized by a mixed-dimensional regime for the right choice of $L_{z}$. Two- and three- dimensional scales are present simultaneously in the flow for $L_{z}=1.67h,3.33h$. In these cases, large scales are two-dimensional while the small ones are three-dimensional, with the transition between regimes happening at some intermediate scale. This transition is strongly dependent on the height of the domain and it can be delayed further in $x$ for increasing values of $L_{z}$.
a)b)c)
@@
Figure 3: Third-order structure functions $S_{3}$ of the longitudinal velocity fluctuations at different distances from the inlet. Panel a shows $S_{3}$ computed at distances $x$ uniformly separated along the jet centerline ($x=20h,30h,\ldots,140h$) for the $2.5$D turbulent planar jet with $L_{z}=3.33h$. We indicate whether the turbulent scales are two- or three- dimensional with colors, similarly to Fig. 2. The plots are shifted vertically for better readability. The inset shows the scale $r$ at which $S_{3}$ changes its sign. Red and blue markers are used for fully $3$D and $2$D structure functions respectively, while purple markers for those showing both regimes. Note that, if $S_{3}$ does not change sign, we mark either the largest three-dimensional scale or the smallest two-dimensional one. Panel b depicts $S_{3}$ for the least constrained ($L_{z}=13.33h$) or $3$D jet, and panel c for the most constrained ($L_{z}=0.83h$) or $2$D jet. $S_{3}$ are plotted in a similar fashion than a, where we computed $S_{3}$ at three distances $x=40h,80h,120h$ at the jet centerline. In these cases, $S_{3}$ has the same sign throughout $x$, thus displaying a single scaling at all distances dependent on the dimensionality of the flow.
@@
To further investigate the transition between regimes, we choose a mixed-dimensional planar jet, in particular $L_{z}=3.33h$, and we calculate repeatedly $S_{3}$ at several distances from the inlet. Results are summarized in Fig. 3a. Close to the inlet the flow is three-dimensional: the effect of the vertical constraint does not hinder the development of the three-dimensional regime. The largest flow scale is smaller than the geometrical constraint, thus not compromising the energy cascade of the three-dimensional flow. We report the presence of a single scaling, $S_{3}\sim-r$, typical of the direct energy cascade. At intermediate distances, $x=\left[40h,100h\right]$, a mixed-dimensional regime appears: both two- and three-dimensional regimes coexist, with three-dimensional turbulence characterizing the small scales and two-dimensional turbulence characterizing instead the large scales. Here, the vertical confinement hinders the development of three-dimensional turbulence at the largest scales, while it has no effect on the three dimensional turbulence at the smallest scales. In this case, we have the simultaneous presence of a direct energy cascade ($S_{3}\sim-r$) at small scales and a direct enstrophy cascade ($S_{3}\sim r^{3}$) at large scales. We report in the inset of Fig. 3a the scale at which the structure function changes sign, i.e., the scale where the flow transitions from $3$D to $2$D, so the largest $3$D scale in the flow. We observe that, as soon as the jet thickness [^1] The jet thickness is defined as the distance from the centerline at which the streamwise velocity equals half of the centerline velocity., which corresponds to the largest $r$, reaches the size of the vertical confinement, the largest scales become two-dimensional. However, some scales between the largest ones and the vertical confinement $L_{z}$ remain three-dimensional, causing the largest $3$D structures to be anisotropic. Moving downstream, the anisotropy of the largest $3$D structures grows till a value of around $3L_{z}$ at $x\approx 60h$, after which the flow becomes more and more two-dimensional. Eventually, we observe that the flow becomes completely two-dimensional at the farthest distances from the inlet, $x\geq 120h$, where the characteristic flow scales are the largest and the vertical confinement impedes the development of three-dimensional flow at all scales. We indeed observe a scaling of $S_{3}$ compatible with the direct enstrophy cascade ($S_{3}\sim r^{3}$). The range of scales observed at each distance from the inlet depends on two factors: the characteristic length scale of the jet (the jet thickness) and the local viscosity. Both of these quantities increase with increasing distance from the inlet of the jet; the jet thickness determines the largest scale in the flow, while the local viscosity is among the factors determining the eventual development of three-dimensional turbulent motions.